What is meant by order and degree of a differential equation. Please explain me with example.

Grade:12

2 Answers

Vijay Mukati
8 years ago
Theorderof adifferential equationis theorderof the highestorderderivative present in theequation. Thedegreeof adifferential equationis the power of the highestorderderivative in theequation.
Example:
d2y/dx2 +(dy/dx) - 3x + 2y = 8
Above differencial equation has order 2 (the highest derivative appearing is the second derivative) and degree 1 (the power of the highest derivative is 1.)
Nicho priyatham
8 years ago
  • Order of a differential equation is defined as the order of the highest order derivative of the dependent variable with respect to the independent variable involved in the given differential equation.
               ex : ( d2y/dx2) + dy/dx +y=0
                       here order is two as the highest order term is   d2y/dx2
  • degree it is the power of  hieght order term
                ex   ( d2y/dx2)4 + dy/dx +y=0
                       
                   here  the hieght order term is  d2y/dxand it is raised to power of 4 degree is 4
                 
  •  Order and degree (if defined) of a differential equation are alway positive integers
  • the dedgree is defined only wn the equation is polynomial in dervative terms it is not define wn deravative term is in exponent function or trignometric or log function
             ex    : (dy/dx) +sin(dy/dx)=0
                             the degree is not define
PLZ APPROVE IF USEFUL and in case of any doubts ask in ans box